### v1

# Self-Supervised Contrastive Learning for Videos using Differentiable Local Alignment

Abstract
Robust frame-wise embeddings are essential to perform video analysis and understanding tasks. We present a self-supervised method for representation learning based on aligning temporal video sequences. Our framework uses a transformer-based encoder to extract frame-level features and leverages them to find the optimal alignment path between video sequences. We introduce the novel Local-Alignment Contrastive (LAC) loss, which combines a differentiable local alignment loss to capture local temporal dependencies with a contrastive loss to enhance discriminative learning. Prior works on video alignment have focused on using global temporal ordering across sequence pairs, whereas our loss encourages identifying the best-scoring subsequence alignment. LAC uses the differentiable Smith-Waterman (SW) affine method, which features a flexible parameterization learned through the training phase, enabling the model to adjust the temporal gap penalty length dynamically. Evaluations show that our learned representations outperform existing state-of-the-art approaches on action recognition tasks.
1 Introduction
Action recognition in videos is crucial in computer vision as it enables the analysis and interpretation of human activity. These actions are inherently sequential, consisting of interconnected events. Typically, tasks in video understanding include video classification [1, 2, 3, 4, 5, 6], which involves mapping a global video representation to an action category, or action detection [7, 8, 9, 10, 11], which aims to identify both the spatial and temporal boundaries of specific actions within a video. However, these methods present several challenges: (i) Most approaches require fine-grained annotations, which are time-consuming and labor-intensive; (ii) Most approaches struggle to capture the complex temporal information and causal relationships that vary across diverse videos. To address these challenges, we explore how representation learning based on the pretext-task of video alignment can improve performance in downstream action recognition tasks such as action phase classification and action progression.
Prior self-supervised or weakly-supervised video alignment approaches [12, 13, 14, 15, 16, 17] train on pairs of videos that describe the same action using either cycle-consistency [13], soft dynamic time warping [15, 14], or sequential contrastive learning [16]. We notice that the pretext task of video alignment shares similarities with bioinformatics sequence alignment for studying DNA, RNA, and protein structures. Since some actions are composed of sub-actions or events that occur in a specific order, video alignment shares similarities with protein alignment, as both involve identifying regions of continuity and discontinuity within sequences. There are two main types of alignment algorithms commonly used in bioinformatics: global and local. The Needleman-Wunsch (NW) algorithm [18], a global alignment method, aligns entire sequences from start to finish. This approach is conceptually similar to Dynamic Time Warping (DTW) [19], which aligns sequences of varying lengths by matching their temporal patterns from start to finish. Differentiable versions of NW and DTW algorithms have been established by incorporating smooth maximum/minimum and argmax/argmin functions [20, 21]. In particular, Soft-DTW [21] has been used in prior video alignment approaches [14, 15, 17]. In contrast, the Smith-Waterman (SW) algorithm [22], which implements local alignment, identifies regions of high similarity within sequences and is adaptable to gaps, insertions, and deletions. This capability to handle subsequences makes local alignment particularly appropriate for managing the complexities of action recognition in realistic settings. Recently, bioinformatics research has introduced a differentiable form of the SW algorithm [23, 24], which we adopt for the first time in the context of representation learning to align video sequences.
Specifically, in this work, we introduce a novel loss function termed Local-Alignment Contrastive (LAC) loss, integrated into our proposed end-to-end framework for the video alignment task, as visualized in Fig. 1. Following prior works [16, 17], our framework uses a variation of convolutional and transformer encoders to extract spatio-temporal features from each frame. Our novel LAC loss enables the comparison of video pairs through a consistent and differentiable local alignment loss. This loss specifically accommodates sequences with learned penalties for opening gaps and extending them, combined with a contrastive loss that effectively separates dissimilar frames. We demonstrate that the learned representations of our approach outperform previous methods on various downstream action recognition tasks, such as action phase classification and action phase progression on the Pouring [12] and PennAction [25] datasets.
Figure 1: We introduce a representation learning approach that aligns video sequences depicting the same processes. Our training objective is to use a novel LAC loss to optimize and learn an element-wise embedding function that supports the alignment process.
2 Related Work
Action Recognition. Over the last decade, researchers have proposed various methods to address video recognition tasks [3, 9, 26, 27, 28]. Video alignment, as a self-supervised pretext task for action recognition, enables the extraction of video representations without requiring labeled data. Previous work has leveraged traditional computer vision techniques to derive these representations. For instance, optical flow [29] computes motion between two frames based on brightness constancy, and the CONDENSATION algorithm [30] employs a probabilistic approach informed by prior motion data. Similarly, Space-Time Interest Points (STIPs) [31] detect salient points in video sequences by adapting image-based point detectors to analyze both spatial and temporal patterns.
Self-Supervised Learning. With the rise of deep learning, the field has shifted towards using Self-Supervised Learning (SSL) methods to learn video representations. Prior works [12, 13, 14, 15, 16] have addressed the problem of video alignment using self-supervised methods. Time Contrastive Network (TCN) [12] utilizes contrastive learning to distinguish frames from different segments while grouping those within the same segment. Temporal Cycle-Consistency (TCC) [13] introduces cycle-consistency loss to identify and match recurring action sequences within or across videos. Learning by Aligning Videos (LAV) [14] uses Soft-DTW [21] and Inverse Difference Moment (IDM) regularization to optimize alignment and ensure balanced frame distribution. Global Temporal Alignment (GTA) [15] implements a modified differentiable DTW with global consistency loss for consistent temporal alignment. Contrastive Action Representation Learning (CARL) [16] introduces Sequence Contrastive Loss (SCL), which minimizes the KL-divergence between augmented views based on timestamp distance. Learning Representation by position PROPagation (LRProp) [17] utilizes a Soft-DTW [21] alignment path and a novel pair-wise position propagation technique. We will use these state-of-the-art methods to evaluate our model’s performance on action recognition tasks.
3 Method
This section introduces our novel Local-Alignment Contrastive (LAC) loss function employed within our framework where video alignment serves as the pretext task, as visualized in Fig. 2. Specifically, we will derive the forward and backward passes of the differentiable local alignment loss and explain how we ensure consistency within our method.
3.1 Notations
Let $V^{1}$ and $V^{2}$ represent a pair of video frames. Following [16], we preprocess the frames via temporal random cropping to generate two cropped videos, each consisting of $T$ frames. We denote these cropped videos as $(V^{1},S^{1})$ and $(V^{2},S^{2})$, where $S\in\mathbb{R}^{T}$ contains the indices of the sampled frames. Next, we apply augmentations, resulting in $\widetilde{V}^{1}$ and $\widetilde{V}^{2}$. Finally, these frames are processed through an encoder to obtain embeddings $Z^{1}$ and $Z^{2}$. Each embedding $Z\in\mathbb{R}^{T\times E}$ represents a sequence with length $T$ and embedding dimension $E$. The objective is to train the encoder so that frames from $V^{1}$ and $V^{2}$ depicting the same action have minimized distances between the corresponding embeddings.
Figure 2: Our framework uses Local Alignment Loss and Contrastive Loss to optimize embeddings generated by an encoder that processes input videos after they have undergone spatio-temporal data augmentation.
3.2 Background
Prior approaches [14, 15] use DTW as the principal method for temporal alignment loss. Given a cost matrix $d_{i,j}$ that measures the cost of matching elements between embeddings $Z^{1}$ and $Z^{2}$, DTW seeks the optimal path that minimizes this cost. The feasible paths must satisfy constraints of matching endpoints, monotonicity, and continuity. Initially, the computational complexity of DTW grows exponentially with the length of the sequences due to the combinatorial explosion of possible alignment paths. However, by using dynamic programming, DTW’s complexity reduced to polynomial time. The DTW algorithm can be formulated as follows:
$D_{i,j}=d_{i,j}+\min\left\{\begin{array}[]{l}D_{i-1,j-1}\\ D_{i-1,j}\\ D_{i,j-1}\\ \end{array}\right.$ (1)
where $D_{i,j}$ represents the cumulative distance up to the point $(i,j)$ in the cost matrix, and $d_{i,j}$ denotes the local distance between elements $Z^{1}_{i}$ and $Z^{2}_{j}$. The non-differentiable nature of the $\min$ operator in the DTW formulation presents challenges for gradient-based optimization techniques. To address this, prior methods [14, 15] utilize Soft-DTW [21], which replaces the discrete $\min$ operator with a differentiable/smoothed version denoted as $\min^{\gamma}$. As the parameter $\gamma$ approaches zero, he behavior of $\min^{\gamma}$ approximates that of the discrete $\min$ operator. This substitution enables the use of gradient-based optimization in DTW while preserving its alignment capabilities.
3.3 Differentiable Local-Alignment
Despite the advantages offered by Soft-DTW [21], it inherently maintains the global alignment characteristic, ensuring continuity throughout the sequence. This attribute of global alignment may prove suboptimal in scenarios where localized alignment strategies, such as those implemented by the SW algorithm [22], are preferable. The SW algorithm [22] supports localized matching, providing enhanced flexibility and increased sensitivity to similarities within subsequences. In response, we propose a differentiable SW local-alignment approach as our method for temporal alignment loss.
We consider a pair of embeddings, $(Z^{1},S^{1})$ and $(Z^{2},S^{2})$, with corresponding sampled indices, tasked with determining the optimal alignment between these embeddings. We begin by calculating a similarity matrix, $S\in\mathbb{R}^{T\times T}$, using the inverted Euclidean distance to yield higher values that indicate increased similarity between the points in $Z^{1}$ and $Z^{2}$. Utilizing dynamic programming, the SW algorithm effectively determines the optimal alignment by maximizing a score derived from pairwise similarities within the matrix. The alignment algorithm can be formally expressed as follows:
$D_{i,j}=S_{i,j}+\max\left\{\begin{array}[]{l}0\\ D_{i-1,j-1}\\ I_{x_{i-1,j-1}}\\ I_{y_{i-1,j-1}}\end{array}\right.\$ (2)
$I_{x_{i,j}}=\max\left\{\begin{array}[]{l}D_{i,j-1}-g_{o_{i,j}}\\ I_{x_{i,j-1}}-g_{e_{i,j}}\end{array}\right.$ (3)
$I_{y_{i,j}}=\max\left\{\begin{array}[]{l}D_{i-1,j}-g_{o_{i,j}}\\ I_{x_{i-1,j}}-g_{o_{i,j}}\\ I_{y_{i-1,j}}-g_{e_{i,j}}\end{array}\right.$ (4)
where the initial values are set to $D(i,0)=D(0,j)=I_{x}(i,0)=I_{x}(0,j)=I_{y}(i,0)=I_{y}(0,j)=-\infty$ for all indices $i,j=1,\ldots,T$. The matrix $D$ described in Eq. 2 acts as the primary scoring matrix, representing the cumulative maximum score at each matrix coordinate $(i,j)$. The matrices $I_{x}$ and $I_{y}$ detailed in Eq. 3 and Eq. 4, represent the scores associated with introducing gaps along the $x$ and $y$ axes. Additionally, $g_{o}$ and $g_{e}$ are used to specify the penalties for opening and extending gaps.
Similar to Soft-DTW [21], our differentiable temporal alignment loss function uses a smoothed version of the discrete $\max$ function, denoted as $\max^{\gamma}$. This smooth approximation is formally defined as:
$\max{}^{\gamma}(u_{1},\ldots,u_{n}):=\gamma\,\log\,\sum_{i=1}^{n}\,e^{u_{i}/\gamma}$ (5)
where $\gamma>0$ is a smoothing parameter that controls the approximation’s fidelity to the original $\max$ function. As $\gamma$ approaches zero, the smoothed $\max^{\gamma}$ function increasingly approximates the behavior of the standard $\max$ function, enabling a differentiable formulation that can be integrated into gradient-based optimization frameworks. The gradient of the smoothed $\max^{\gamma}$ with respect to each input $u_{i}$ is calculated as follows:
$\frac{\partial\max{}^{\gamma}(\{u_{1},\ldots,u_{n}\})}{\partial u_{i}}$ $=\exp\left(\frac{u_{i}-\max{}^{\gamma}(\{u_{1},\ldots,u_{n}\})}{\gamma}\right)$ (6)
This expression aligns with the softmax function. By differentiating the smoothed maximum score with respect to each input, our model directly utilizes derivative-sensitive parameters such as the similarity score ($S$), gap open penalty ($g^{o}$), and gap extend penalty ($g^{e}$). As a result, we define our differentiable local alignment loss as $\mathcal{L}_{sw_{ij}}=\text{SW}(\text{sim}(Z^{i},Z^{j}))$. The (sim) function calculates the similarity between two embeddings and generates a similarity matrix $S$. Unlike prior differentiable SW implementations [23, 24] that depend on supervised learning, our approach integrates the differentiable SW algorithm within a self-supervised framework.
3.4 Final Loss
Following the approach in [16], our contrastive loss is modeled after the NT-Xent loss from SimCLR [32]. This loss function calculates absolute pairwise distances between embeddings $(Z^{1},S^{1})$ and $(Z^{2},S^{2})$, subsequently forming a Gaussian-weighted positive label distribution. This distribution is then contrasted against normalized similarity logits using the Kullback-Leibler divergence. The contrastive loss is mathematically expressed as:
$\mathcal{L}_{c}=-\frac{1}{T}\sum_{i,j=1}^{T}\left(\frac{G(s_{i}^{1}-s_{j}^{2})}{\sum_{k=1}^{T}G(s_{i}^{1}-s_{k}^{2})}\log\frac{\exp(\text{sim}(z_{i}^{1},z_{j}^{2})/\tau)}{\sum_{k=1}^{T}\exp(\text{sim}(z_{i}^{1},z_{k}^{2})/\tau)}\right)$ (7)
[…]
$\mathcal{L}_{l}$ $=-\frac{1}{T}\sum_{i,j=1}^{T}\left(\,\frac{G(s_{i}^{1}-s_{j}^{2})}{\sum_{k=1}^{T}G(s_{i}^{1}-s_{k}^{2})}\log\left(\frac{\exp(\text{L}_{ij})}{\sum_{k=1}^{T}\exp^{(}\text{L}_{ik})}\right)\right),$
$\text{where}\,\ \tilde{D}_{12,ij}$ $=\frac{\exp(D_{12,ij}/\tau)}{\sum_{k=1}^{T}\exp(D_{12,ik}/\tau)}\,,\,\ \tilde{D}_{21,ij}=\frac{\exp(D_{21,ij}/\tau)}{\sum_{k=1}^{T}\exp(D_{21,ik}/\tau)}\,,\,\ L=\tilde{D}_{12}\cdot\tilde{D}_{21}^{T}$
Our proposed loss function, referred as Local-Alignment Contrastive (LAC), integrates contrastive and local alignment losses. It is formally defined as follows:
$\mathcal{L}=\mathcal{L}_{c}+\alpha\cdot(\mathcal{L}_{l}+\beta\cdot(\mathcal{L}_{sw12}+\mathcal{L}_{sw21}))$ (8)
where $\alpha$ and $\beta$ are weights used to balance the components of the loss, set to 0.01 and 1, respectively.
4 Experiments
Datasets. We evaluate the performance of the LAC model on two datasets using a variety of evaluation metrics. The Pouring dataset [12], which focuses on the action of pouring liquids, includes 70 training and 14 testing videos. The PennAction dataset [25], featuring 13 human actions, contains 1140 training and 966 testing videos. We utilize the key events and phases for the videos in both datasets as proposed by TCC [13].
Implementation Details. Our encoder, $f:\mathbb{R}^{TxCxWxH}\to Z$, maps video inputs $V$ into an embedding space $Z$. We use a ResNet50-v2 [33] as our backbone to extract features from the Conv4c layer with an output size of $51210x10x$ 512. These features are then processed through adaptive max pooling, followed by two fully connected layers with ReLU activation. A subsequent linear layer projects the features into a 256-dimensional space. To enhance the model’s capacity to capture long-range dependencies, we integrate sine-cosine positional encoding and employ a two-layer Transformer encoder. To improve the model’s ability to capture long-range dependencies, we incorporate sine-cosine positional encoding and apply a two-layer Transformer encoder. The final embedding layer reduces the dimensionality to 128 for the frame-wise representations.
Method Class AP@K Progress $\tau$
10 50 100 K=5 K=10 K=15
TCN [12] 80.32 81.44 83.56 76.26 76.71 77.26 82.30 83.51
TCC [13] 86.60 86.78 86.86 81.84 80.94 81.69 83.36 85.26
LAV [14] 89.77 90.35 91.77 87.48 88.36 88.40 85.20 88.75
GTA [15] 89.34 90.20 90.22 87.79 87.48 87.82 88.67 92.47
SCL [16] 92.76 92.80 93.05 88.75 88.51 88.97 91.26 98.20
LRPROP [17] 92.70 94.44 94.36 92.41 90.33 90.86 94.09 99.46
LAC 95.87 95.78 95.16 92.76 91.07 91.37 94.24 97.50
Table 1: Performance comparison of state-of-the-art methods on the Pouring Dataset [12]
Evaluation Metrics. Following related work [13, 14, 15, 16, 17], we evaluate our model using the following metrics: (i) Phase Classification, which assesses the accuracy of action phase predictions by training an SVM classifier on our embeddings; (ii) Phase Progression, which evaluates how accurately our embeddings predict action progress using a linear regression model’s average R-squared value, based on normalized timestamp differences; (iii) Average Precision@K (AP@K), which evaluates fine-grained frame retrieval accuracy by calculating the proportion of correctly matched phase labels within the K closest frames; (iv) Kendall’s Tau, which quantifies the temporal alignment between sequences by comparing the ratio of concordant to discordant frame pairs.
4.1 Results
We apply the same four metrics to the Pouring dataset [12]. For the PennAction dataset [25], we evaluate each metric across all action categories and report the average results. To ensure a fair comparison, we replicated the evaluations of previous approaches using the GitHub repositories 111github.com/google-research/google-research/tcc, github.com/trquhuytin/LAV-CVPR21,
github.com/hadjisma/VideoAlignment, github.com/minghchen/CARL_code
provided by the original authors. Each model was trained using its respective pre-trained backbone. An exception was made for LRPROP, as their GitHub repository is not available.
Results on Pouring Dataset. Table 1 presents a comparison of our method’s performance against state-of-the-art approaches on the Pouring dataset. Bold and underlined text denote the best and second-best results. Notably, it achieves a +3.11% improvement on Phase Classification using only 10% of the labels. Additionally, our model excels in AP@K and Progress metrics. However, we observe lower performance in Kendall’s Tau. We hypothesize this is due to how the SW’s algorithm encourages skipping unnecessary segments of the sequence. The introduction of gap open and extend penalties could disrupt the continuity needed for high Kendall’s Tau scores.
Method Class AP@K Progress $\tau$
10 50 100 K=5 K=10 K=15
TCN [12] 69.73 70.26 70.01 60.92 61.57 61.52 76.37 63.72
TCC [13] 86.60 86.78 86.86 81.84 80.94 81.69 83.36 85.26
LAV [14] 88.51 88.72 88.97 73.47 73.13 74.27 92.52 93.06
GTA [15] 84.21 84.68 85.28 71.72 72.17 71.52 90.51 83.35
SCL [16] 87.85 87.52 88.15 91.70 90.61 90.58 92.89 98.14
LRPROP [17] 91.90 92.96 93.25 92.46 92.2 92.03 93.03 99.09
LAC 95.57 93.79 93.40 93.87 93.41 92.65 94.21 94.10
Table 2: Performance comparison of state-of-the-art methods on the PennAction Dataset [25]
Refer to caption Refer to caption
Figure 3: Similarity matrix (left) shows video alignment using an optimal path and the respective video aligned frame-by-frame (right).
Results on PennAction Dataset. Table 2 compares the performance of our method with state-of-the-art approaches on the PennAction dataset. Bold and underlined text denote the best and second-best results. LAC consistently outperforms previous methods across most metrics, with the exception of Kendall’s Tau. Notably, the improvement in AP@K is more pronounced on PennAction than on Pouring, likely due to the fewer number of frames in PennAction dataset, which may further impact alignment performance.
Refer to caption Refer to caption
Figure 4: Fine-grained frame retrieval for Pouring (left) and PennAction (right) achieved using nearest neighbors within our embedding space.
Qualitative analysis of results. Figure 3 illustrates the alignment process, with the optimal path depicted on the left and the frame-by-frame alignment on the right. This visualization demonstrates the synchronization of the two videos despite differences in their temporal progression and duration. Such alignment visualizations enable in-depth analysis of deviations or inefficiencies within specific actions in video understanding. Additionally, Figure 4 showcases our embedding-based retrieval system’s ability to accurately identify and retrieve frames corresponding to specific action sequences across different videos.
4.2 Ablation Study
This section presents multiple experiments on the Pouring dataset that analyze the different components of our framework.
Architecture Class Progress $\tau$
[…]
ResNet-50 + Tranformer (w/o pretrained weights) 90.06 89.32 94.87
ResNet-50 + Transformer 95.16 94.24 97.50
Table 3: Different Encoder Architecture on Model Performance.
$\gamma$ Class Progress $\tau$
[…]
0.7 94.87 92.82 97.29
0.8 95.16 94.24 97.50
0.9 93.93 92.02 94.91
Table 4: Different $\gamma$ Values on Model Performance
Loss $\mathcal{L}$ Class Progress $\tau$
[…]
$\mathcal{L}_{c}+\alpha\cdot\mathcal{L}_{l}$ 93.71 92.16 95.01
$\mathcal{L}_{c}+\alpha\cdot(\mathcal{L}_{l}+\beta\cdot(\mathcal{L}_{sw12}+\mathcal{L}_{sw21}))$ 95.16 94.24 97.50
Table 5: Different LAC Loss Formulations on Model Performance
Network Architecture. Table 3 demonstrates that our performance evaluation across various network architectures highlights the superiority of the ResNet-50 model with pretrained weights combined with a Transformer.
Hyperparameter of LAC Loss. Table 5 presents the optimal performance results obtained at the smoothing parameter $\gamma=0.8$. Table 5 further demonstrates that integrating local alignment loss into the total loss significantly enhances performance and achieves the highest scores across all metrics compared to using contrastive loss alone.
5 Conclusion
The paper introduces a novel approach to representation learning through a Local-Alignment Contrastive (LAC) loss that integrates a differentiable local alignment loss with a contrastive loss, all within a self-supervised framework. The method employs a differentiable affine Smith-Waterman algorithm to enable temporal alignment that dynamically adjusts to variations in action sequences. This approach is distinct in its focus on capturing local temporal dependencies and enhancing the discriminative learning of video embeddings, accommodating differences in action lengths and sequences. Experimental results on the Pouring and PennAction datasets showcase the method’s superior performance over existing state-of-the-art approaches.